Walter and John leave in their cars at the same time. Walter drives 252 miles. John drives 10 miles per hours faster than Walter and travels 312 miles. If Walter and John arrive at their destinations at the same time, determine how fast Walter is driving.
Answer by mananth(16949)

Walter x mph
John 10 faster x + 10 mph

Walter 252 miles
John 312 miles
Walter time = John time
t=d/r
252 / x = 312 /( x + 10 )
cross multiply
252 ( x + 10 ) = 312 x
252 x + 2520 = 312 x
252 x - 312 x = -2520
-60 x = -2520
/ -60
x = 42 mph
Walter speed = 42 mph
John speed = 52 mph